sierra

Sierra is making a quilt and she has determined she needs 1012 square inches of green fabric and 809 square inches of burgundy. How many square yards of each material will she need? Round your answers up to the nearest quarter yard.
Answer #1

In this problem, we must convert square inches to square yards.


Now, 

the total quantity of green fabric needed in square inches = 1012

the total quantity of burgundy needed in square inches = 809


As we know, 1 yard = 36 inches

=>  1 square yard = 36*36 square inches = 1296 square inches

=>  1 square inch = 1/1296 square yard


So,

The total quantity of green fabric needed in square yards = 1012/1296 = 0.78


The nearest quarter yard is 0.75, so the answer is 0.75 square yards


The total quantity of burgundy needed in square yards = 809/1296 = 0.624


The nearest quarter yard is 0.50, so the answer is 0.50 square yards.


answered by: gavin